# InstructIR: High-Quality Image Restoration Following Human Instructions

### v4

@@ Abstract
Figure 1: Given an image and a prompt for how to improve that image, our all-in-one restoration model corrects the image considering the human instruction. InstructIR, can tackle various types and levels of degradation, and it is able to generalize in some real-world scenarios.
@@ 1 Introduction
Some methods focus on specific degradations, for instance reducing noise (denoising) [102, 103, 64], removing blur (deblurring) [58, 105], or clearing haze (dehazing) [66, 16]. Such methods are effective for their specific task, yet they do not generalize well to other types of degradation. Other approaches use a general neural network for diverse tasks [74, 95, 82, 10], yet training the neural network for each specific task independently. Since using a separate model for each possible degradation is resource-intensive, recent approaches propose All-in-One restoration models [42, 61, 60, 100]. These approaches use a single deep blind restoration model considering multiple degradation types and levels. Contemporary works such as PromptIR [61] or ProRes [49] utilize a unified model for blind image restoration using learned guidance vectors, also known as “prompt embeddings”, in contrast to raw user prompts in text form, which we use in this work.
@@
We propose the first approach that utilizes real human-written instructions to solve inverse problems and image restoration. Our comprehensive experiments demonstrate the potential of using text guidance for image restoration and enhancement by achieving state-of-the-art performance on various image restoration tasks, including image denoising, deraining, deblurring, dehazing and low-light image enhancement. Our model, InstructIR, is able to generalize to restoring images using arbitrary human-written instructions. Moreover, our single all-in-one model covers more tasks than many previous works. We show diverse restoration samples of our method in Figure 1.
@@ 2 Related Work
Recent deep learning methods [16, 64, 58, 44, 95, 74] have shown consistently better results compared to traditional techniques for blind image restoration [29, 18, 73, 35, 54, 37]. The proposed neural networks are based on convolutional neural networks (CNNs) and Transformers [76] (or related attention mechanisms).
We focus on general-purpose restoration models [44, 95, 82, 10]. For example, SwinIR [44], MAXIM [74] and Uformer [82]. These models can be trained -independently- for diverse tasks such as denoising, deraining or deblurring. Their ability to capture local and global feature interactions, and enhance them, allows the models to achieve great performance consistently across different tasks. For instance, Restormer [95] uses non-local blocks [79] to capture complex features across the image.
NAFNet [10] is an efficient alternative to complex transformer-based methods. The model uses simplified channel attention, and gating as an alternative to non-linear activations. The builing block (NAFBlock) follows a simple meta-former [92] architecture with efficient inverted residual blocks [31]. In this work, we build our InstructIR model using NAFNet as backbone, due to its efficient and simple design, and high performance in several restoration tasks.
@@
Single degradation (or single task) restoration methods are well-studied, however, their real-world applications are limited due to the required resources i.e. allocating different models, and select the adequate model on demand. Moreover, images rarely present a single degradation, for instance noise and blur are almost ubiquitous in any image capture.
@@
We use as reference AirNet [42], IDR [100] and ADMS [60]. We also consider the contemporary work PromptIR [61]. The methods use different techniques to guide the blind model in the restoration process. For instance, an auxiliary model for degradation classification [42, 60], or multi-dimensional guidance vectors (also known as “prompts”) [61, 49] that help the model to discriminate the different types of degradation in the image.
Despite it is not the focus of this work, we acknowledge that real-world image super-resolution is a related problem [44, 106, 48, 12], since the models aim to solve an inverse problem considering multiple degradations (blur, noise and downsampling).
@@
In the recent years, multiple methods have been proposed for text-to-image generation and text-based image editing works [4, 53, 70, 34, 30]. These models use text prompts to describe images or actions, and powerful diffusion-based models for generating the corresponding images. Our main reference is InstructPix2Pix [4], this method enables editing from instructions that tell the model what action to perform, as opposed to text labels, captions or descriptions of the input or output images. Therefore, the user can transmit what to do in natural written text, without requiring to provide further image descriptions or sample reference images.
@@ 3.1 Generating Prompts for Training
Why instructions? Inspired by InstructPix2Pix [4], we adopt human written instructions as the mechanism of control for our model. There is no need for the user to provide additional information, such as example clean images, or descriptions of the visual content. Instructions offer a clear and expressive way to interact, enabling users to pinpoint the unpleasant effects (degradations) in the images.
Handling free-form user prompts rather than fixed degradation-specific prompts increases the usability of our model for laypeople who lack domain expertise. We thus want our model to be capable of understanding diverse prompts posed by users “in-the-wild” e.g. kids, adults, or photographers. To this end, we use a large language model (i.e., GPT-4) to create diverse requests that might be asked by users for the different degradations types. We then filter those generated prompts to remove ambiguous or unclear prompts (e.g., “Make the image cleaner”, “improve this image”). Our final instructions set contains over 10000 different prompts in total, for 7 different tasks. We display some examples in Table 1. As we show in Figure 2 the prompts are sampled randomly depending on the input degradation.
Table 1: Examples of our curated GPT4-generated user prompts with varying language and domain expertise.
@@ 3.2 Text Encoder
A text encoder maps the user prompt to a fixed-size vector representation (a text embedding). The related methods for text-based image generation [67] and manipulation [4, 3] often use the text encoder of a CLIP model [62] to encode user prompts as CLIP excels in visual prompts. However, user prompts for degradation contain, in general, little to no visual content (e.g. the use describes the degradation, not the image itself), therefore, the large CLIP encoders (with over 60 million parameters) are not suitable – especially if we require efficiency.
We opt, instead, to use a pure text-based sentence encoder [63], that is, a model trained to encode sentences in a semantically meaningful embedding space. Sentence encoders –pre-trained with millions of examples– are compact and fast in comparison to CLIP, while being able to encode the semantics of diverse user prompts. For instance, we use the BGE-micro-v2 sentence transformer.
@@
Figure 3 shows that while the text encoder is capable out-of-the-box to cluster instructions to some extent (Figure 3(a)), our trained projection yields greatly improved clusters (Figure 3(b)). We distinguish clearly the clusters for deraining, denoising, dehazing, deblurring, and low-light image enhancement. The instructions for such tasks or degradations are very characteristic. Furthermore, we can appreciate that “super-res” and “enhancement” tasks are quite spread and between the previous ones, which matches the language logic. For instance “add details to this image” could be used for enhancement, deblurring or denosising. In our experiments, $d_{t}\!=\!384$, $d_{v}\!=\!256$ and $\mathbf{W}$ is a linear layer. The representation $\mathbf{e}$ from the text encoder is shared across the blocks, and each block has a trainable projection $\mathbf{W}$.
@@ 3.3 InstructIR
Figure 4: Instruction Condition Block (ICB) using an approximation of task routing [71] for many-tasks learning. See Eq. 2.
Our method InstructIR consists of an image model and a text encoder. We introduced our text encoder in Sec. 3.2. We use NAFNet [10] as the image model, an efficient image restoration model that follows a U-Net architecture [68]. To successfully learn multiple tasks using a single model, we use task routing techniques. Our framework for training and evaluating the model is illustrated in Figure 2.
@@
The key aspect of InstructIR is the integration of the encoded instruction as a mechanism of control for the image model. Inspired in task routing for many-task learning [69, 71, 14], we propose an “Instruction Condition Block” (ICB) to enable task-specific transformations within the model. Conventional task routing [71] applies task-specific binary masks to the channel features. Since our model does not know a-priori the degradation, we cannot use this technique directly.
@@
where the mask $\mathbf{m}_{c}=\sigma(\mathbf{W_{c}}\cdot\mathbf{e})$ is produced using a linear layer -activated using the Sigmoid function- to produce a set of weights depending on the text embedding $\mathbf{e}$. Thus, we obtain a $c$-dimensional per-channel (soft-)binary mask $\mathbf{m}_{c}$. As [71], task routing is applied as the channel-wise multiplication $\odot$ for masking features depending on the task. The conditioned features are further enhanced using a NAFBlock [10] ($\mathrm{Block}$). We illustrate our task-routing ICB block in Figure 4. We use “regular” NAFBlocks [10], followed by ICBs to condition the features, at both encoder and decoder blocks. The formulation is $F^{l+1}\!=\!\mathrm{ICB}(\mathrm{Block}(F^{l}))$ where $l$ is the layer. Although we do not condition explicitly the filters of the neural network, as in [71], the mask allows the model to select the most relevant channels depending on the image information and the instruction. Note that this formulation enables differentiable feature masking, and certain interpretability i.e. the features with high weights contribute the most to the restoration process. Indirectly, this also enforces to learn diverse filters and reduce sparsity [71, 14].
@@ 4 Experimental Results
We provide extensive qualitative results using benchmark images in Figures 1920 and 21. We also evaluate our model on 9 well-known benchmarks for different image restoration tasks: image denoising, deblurring, deraining, dehazing, and image enhancement. We present extensive quantitative results in Table 2. Our single model successfully restores images considering different degradation types and levels. We provide additional results and ablation studies in the supplementary material.
@@
Our InstructIR model is end-to-end trainable. The image model does not require pre-training, yet we use a pre-trained sentence encoder as language model.
@@
We use NAFNet [10] as image model. The architecture consists of a 4-level encoder-decoder, with varying numbers of blocks at each level, specifically [2, 2, 4, 8] for the encoder, and [2, 2, 2, 2] for the decoder, from the level-1 to level-4 respectively. Between the encoder and decoder we use 4 middle blocks to enhance further the features. The decoder implements addition instead of concatenation for the skip connections.
We use the Instruction Condition Block (ICB) for task-routing [71] only in the encoder and decoder.
The model is optimized using the $\mathcal{L}_{1}$ loss between the ground-truth clean image and the restored one. Additionally we use the cross-entropy loss $\mathcal{L}_{ce}$ for the intent classification head of the text encoder. We train use a batch size of 32 and AdamW [36] optimizer with learning rate $5e^{-4}$ for 500 epochs (approximately 1 day using a single NVIDIA A100). We also use cosine annealing learning rate decay. During training, we utilize cropped patches of size $256\times 256$ as input, and we use random horizontal and vertical flips as augmentations. Since our model uses as input instruction-image pairs, given an image, and knowing its degradation, we randomly sample instructions from our prompt dataset ($>\!10$K samples). Our image model has only 16M parameters, and the learned text projection is just $100$k parameters (the language model is 17M parameters), thus, our model can be trained easily on standard GPUs such as NVIDIA RTX 2080Ti or 3090Ti in a couple of days. Furthermore, the inference process also fits in low-computation budgets.
@@ 4.2 Datasets and Benchmarks
Following previous works [42, 100, 61], we prepare the datasets for different restoration tasks.
@@
We use the GoPro dataset for motion deblurring [57] which consist of 2103 images for training, and $1$ for testing.
Low-light Image Enhancement.
We use the LOL [83] dataset (v1), and we adopt its official split of 485 training images, and 15 testing images.
Image Enhancement.
Extending previous works, we also study photo-realistic image enhancement using the MIT5K dataset [5]. We use 1000 images for training, and the standard split of 500 images for testing (as in [74]).
Table 2: Quantitative results on five restoration tasks (5D) with state-of-the-art general image restoration and all-in-one methods. We highlight the reference model without text (image only), the best overall results, and the second best results. We also present the ablation study of our multi-task variants (from 5 to 7 tasks — 5D, 6D, 7D). This table is based on Zhang et al. IDR [100] (CVPR ’23).
@@
Finally, as previous works [42, 100, 61], we combine all the aforementioned training datasets, and we train our unified model for all-in-one restoration.
4.3 Multiple Degradation Results
We define two initial setups for multi-task restoration:
In Table 2, we show the performance of 5D models. Following Zhang et al. [100], we compare InstructIR with several state-of-the-art methods for general image restoration [95, 10, 9, 44, 93], and all-in-one image restoration methods [100, 42, 75, 21, 45]. We can observe that our simple image model (just 16M parameters) can tackle successfully at least five different tasks thanks to the instruction-based guidance, and achieves the most competitive results.
In Table 4 we can appreciate a similar behaviour, when the number of tasks is just three (3D), our model improves further in terms of reconstruction performance.
Table 3: Ablation study on the sensitivity of instructions. We report PSNR/SSIM metrics for each task using our 5D base model. We repeat the evaluation on each testset 10 times, each time we sample different prompts for each image, and we report the average results. The “Real Users $\dagger$” in this study are amateur photographers, thus, the instructions were very precise.
Language Level Deraining Denoising Deblurring LOL
Basic & Precise 36.84/0.973 31.40/0.887 29.47/0.887 23.00/0.836
Basic & Ambiguous 36.24/0.970 31.35/0.887 29.21/0.885 21.85/0.827
Test Instructions.
InstructIR requires as input the degraded image and the human-written instruction. Therefore, we also prepare a testset of prompts i.e. instruction-image test pairs. The performance of InstructIR depends on the ambiguity and precision of the instruction. We provide the ablation study in Table 3. InstructIR is quite robust to more/less detailed instructions. However, it is still limited with ambiguous instructions such as “enhance this image”. We show diverse instructions in the following Figures.
Table 4: Comparisons of all-in-one restoration models for three restoration tasks (3D). We also show an ablation study for image denoising -the fundamental inverse problem- considering different noise levels. We report PSNR/SSIM metrics. Table based on [61].
@@ 4.3 Multiple Degradation Results
Rain, Blur and Noise “Correct the noise” “Remove the rain” “Increase resolution” “Enhance the photo”
Figure 5: Selective task. InstructIR can remove particular degradations or perform different transformations depending on the human instructions. This is a novel feature in image restoration, and it is possible thanks to the novel integration of textual descriptions.
5 Multi-Task Ablation Study
Table 5: Image Enhancement performance on MIT5K [5, 96].

### v5

# %****␣eccv_paper.tex␣Line␣100␣****https://github.com/mv-lab/InstructIR (500 ✰)

@@ Abstract
Figure 1: Given an image and a prompt for how to improve that image, our all-in-one restoration model corrects the image considering the human instruction. InstructIR, can tackle various types and levels of degradation, and it is able to generalize in some real-world scenarios (last three images, from left to right).
@@ 1 Introduction
Some methods focus on specific degradations, for instance reducing noise (denoising) [103, 105, 65], removing blur (deblurring) [59, 107], or clearing haze (dehazing) [67, 16]. Such methods are effective for their specific task, yet they do not generalize well to other types of degradation. Other approaches use a general neural network for diverse tasks [75, 95, 83, 9], yet training the neural network for each specific task independently. Since using a separate model for each possible degradation is resource-intensive, recent approaches propose All-in-One restoration models [43, 62, 61, 102]. These approaches use a single deep blind restoration model considering multiple degradation types and levels. Contemporary works such as PromptIR [62] or ProRes [50] utilize a unified model for blind image restoration using learned guidance vectors, also known as “prompt embeddings", in contrast to raw user prompts in text form, which we use in this work.
@@
We propose the first approach that utilizes real human-written instructions to solve multi-task image restoration. Our comprehensive experiments demonstrate the potential of using text guidance for image restoration and enhancement by achieving state-of-the-art performance on various image restoration tasks, including image denoising, deraining, deblurring, dehazing, and low-light image enhancement. Our model, InstructIR, is able to generalize to restoring images using complex human-written instructions. Moreover, our single all-in-one model covers more tasks than many previous works. We show diverse restoration samples of our method in Figure 1.
@@ 2 Related Work
Recent deep learning methods [16, 65, 59, 45, 95, 75] have shown consistently better results compared to traditional techniques for blind image restoration [30, 18, 74, 36, 55, 38]. The proposed neural networks are based on convolutional neural networks (CNNs) and Transformers [77] (or related attention mechanisms). We focus on general-purpose restoration models [45, 95, 83, 9]. For example, SwinIR [45], MAXIM [75] and Uformer [83]. These models can be trained -independently- for diverse tasks such as denoising, deraining or deblurring. Their ability to capture local and global feature interactions, and enhance them, allows the models to achieve great performance consistently across different tasks. For instance, Restormer [95] uses non-local blocks [80] to capture complex features across the image.
NAFNet [9] is an efficient alternative to complex transformer-based methods. The model uses simplified channel attention, and gating as an alternative to non-linear activations. The building block (NAFBlock) follows a simple meta-former [94] architecture with efficient inverted residual blocks [32]. In this work, we build our InstructIR model using NAFNet as backbone, due to its efficient and simple design, and high performance in several restoration tasks.
@@
Single degradation (or single task) restoration methods are well-studied, however, their real-world applications are limited due to the required resources i.e. allocating different models, and selecting the adequate model on demand. Moreover, images rarely present a single degradation, for instance, noise and blur are almost ubiquitous in any image capture.
@@
We use as reference AirNet [43], IDR [102] and ADMS [61]. We also consider the contemporary work PromptIR [62]. The methods use different techniques to guide the blind model in the restoration process. For instance, an auxiliary model for degradation classification [43, 61], or multi-dimensional guidance vectors (also known as “prompts") [62, 50, 49] that help the model to discriminate the different types of degradation in the image.
@@
In recent years, multiple methods have been proposed for text-to-image generation and text-based image editing works [4, 54, 71, 35, 31]. These models use text prompts to describe images or actions, and powerful diffusion-based models for generating the corresponding images. Our main reference is InstructPix2Pix [4], this method enables editing from instructions that tell the model what action to perform, as opposed to text labels, captions or descriptions of the input or output images. Therefore, the user can transmit what to do in natural written text, without requiring to provide further image descriptions or sample reference images.
@@ 3.1 Generating Prompts for Training
Why instructions? Inspired by InstructPix2Pix [4], we adopt human written instructions as the mechanism of control for our model. There is no need for the user to provide additional information, such as example clean images, or descriptions of the visual content. Instructions offer a clear and expressive way to interact, enabling users to pinpoint the unpleasant effects (degradations) in the images. We also consider the language complexity, from ambiguous instructions (e.g. “fix my image”) to precise instructions (e.g. “remove the noise”).
Table 1: Examples of our curated GPT4-generated and real user prompts with varying language and domain expertise.
@@
General Fix my image please
make the image look better
Handling free-form user prompts rather than fixed degradation-specific prompts increases the usability of our model for laypeople who lack domain expertise. We thus want our model to be capable of understanding diverse prompts posed by users “in-the-wild" e.g. kids, adults, or photographers. To this end, we use a large language model (i.e., GPT-4) to create diverse requests that might be asked by users for the different degradations types. We then filter those generated prompts to remove ambiguous or unclear prompts (e.g., “Make the image cleaner”, “improve this image”). Our final instructions set contains over 10000 different prompts in total, for 7 different tasks. We display some examples in Table 1. As we show in Figure 2 the prompts are sampled randomly depending on the input degradation.
@@ 3.2 Text Encoder
A text encoder maps the user prompt to a fixed-size vector representation (a text embedding). The related methods for text-based image generation [68] and manipulation [4, 3] often use the text encoder of a CLIP model [63] to encode user prompts as CLIP excels in visual prompts. However, user prompts for degradation contain, in general, little to no visual content (e.g. the use describes the degradation, not the image itself). We opt, instead, to use a pure text-based sentence encoder [64], that is, a smaller model trained to encode sentences in a semantically meaningful embedding space. Sentence encoders – pre-trained with millions of examples – are compact and fast in comparison to CLIP, while being able to encode the semantics of diverse user prompts. For instance, we use the BGE-micro-v2 sentence transformer. We compare text encoders in the supplementary material.
@@
Figure 3 shows that while the text encoder is capable out-of-the-box to cluster instructions to some extent (Figure 3(a)), our trained projection yields greatly improved clusters (Figure 3(b)). We distinguish clearly the clusters for deraining, denoising, dehazing, deblurring, and low-light image enhancement. The instructions for such tasks or degradations are very characteristic. Furthermore, we can appreciate that “super-res" and “enhancement" tasks are quite spread and between the previous ones, which matches the language logic. For instance “add details to this image" could be used for enhancement, deblurring, or denoising. In our experiments, $d_{t}\!=\!384$, $d_{v}\!=\!256$ and $\mathbf{W}$ is a linear layer. The representation $\mathbf{e}$ from the text encoder is shared across the blocks, and each block has a trainable projection $\mathbf{W}$.
@@ 3.3 InstructIR
Our method InstructIR consists of an image model and a text encoder. We introduced our text encoder in Sec. 3.2. We use NAFNet [9] as the image model, an efficient image restoration model that follows a U-Net architecture [69]. To successfully learn multiple tasks using a single model, we use task routing techniques. Our framework for training and evaluation is illustrated in Figure 2.
Figure 4: Instruction Condition Block (ICB) using an approximation of task routing [72] for many-tasks learning (See Eq. 2). This mechanism allows the neural network to select and prioritize specific features depending on the instruction, similarly to a Mixture of Experts (MoE).
@@
The key aspect of InstructIR is the integration of the encoded instruction as a mechanism of control for the image model. Inspired in task routing for many-task learning [70, 72, 14], we propose an “Instruction Condition Block" (ICB) to enable task-specific transformations within the model. Conventional task routing [72] applies task-specific binary masks to the channel features. Since our model does not know a-priori the degradation, we cannot use this technique directly.
@@
where the mask $\mathbf{m}_{c}=\sigma(\mathbf{W_{c}}\cdot\mathbf{e})$ is produced using a linear layer $\mathbf{W_{c}}$ – activated using the Sigmoid function – to produce a set of weights depending on the text embedding $\mathbf{e}$. Thus, we obtain a $c$-dimensional per-channel (soft-)binary mask $\mathbf{m}_{c}$. As [72, 29], task routing is applied as the channel-wise multiplication $\odot$ for masking features depending on the task. The conditioned features are further enhanced using a convolutional NAFBlock [9] ($\mathrm{Block}$). We illustrate our task-routing ICB block in Figure 4. We use “regular” NAFBlocks [9], followed by ICBs to condition the features, at both encoder and decoder blocks. The formulation is $F^{l+1}\!=\!\mathrm{ICB}(\mathrm{Block}(F^{l}))$ where $l$ is the layer. Although we do not condition explicitly the filters of the neural network, as in [72], the mask allows the model to select the most relevant channels depending on the image information and the instruction. Note that this formulation enables differentiable feature masking, and certain interpretability i.e. the features with high weights contribute the most to the restoration process. Indirectly, this also enforces to learn diverse filters and reduce sparsity [72, 14].
@@ 4 Experimental Results
We evaluate our model on 9 well-known benchmarks for different image restoration tasks: image denoising, deblurring, deraining, dehazing, real low-light enhancement, and photo-realistic image enhancement. We present extensive quantitative results in Table 2 and Table 3. We provide extensive comparisons with other all-in-one methods as well as task-specific methods. Our single model successfully restores images considering different degradation types and levels.
@@
Our InstructIR model is end-to-end trainable. The image model does not require pre-training but we use a pre-trained sentence encoder as language model.
@@
We use NAFNet [9] as the image model backbone. The architecture consists of a 4-level encoder-decoder, with varying numbers of blocks at each level, specifically [2, 2, 4, 8] for the encoder, and [2, 2, 2, 2] for the decoder, from the level-1 to level-4 respectively. Between the encoder and decoder we use 4 middle blocks to enhance further the features. The decoder implements addition instead of concatenation for the skip connections. We use the Instruction Condition Block (ICB) for task-routing [72] only in the encoder and decoder.
The model is optimized using the $\mathcal{L}_{1}$ loss between the ground-truth clean image and the restored one. Additionally, we use the cross-entropy loss $\mathcal{L}_{ce}$ for the intent classification head of the text encoder. We train using a batch size of 32 and AdamW [37] optimizer with learning rate $5e^{-4}$ for 500 epochs (approximately 1 day using a single NVIDIA A100). We also use cosine annealing learning rate decay. During training, we utilize cropped patches of size $256\times 256$ as input, and we use random horizontal and vertical flips as augmentations. Since our model uses as input instruction-image pairs, given an image, and knowing its degradation, we randomly sample instructions from our prompt dataset ($>\!10$K samples). Our image model has only 16M parameters, and the learned text projection is just $100$k parameters (the language model is 17M parameters), thus, our model can be trained easily on standard GPUs, furthermore, the inference process also fits in low-computation budgets (e.g. Google Colab T4 16Gb GPU).
@@ 4.2 Datasets and Benchmarks
Following previous works [43, 102, 62], we prepare the datasets for different restoration tasks, including real and synthetic datasets.
@@
We use the GoPro dataset for motion deblurring [58] which consists of 2103 images for training, and $1$ for testing.
Real-world Low-light Image Enhancement.
We use the LOL [84] dataset (v1), which contains real-case low/normal-light image pairs. We adopt its official split of 485 training images and 15 testing images.
Real-world Image Enhancement.
Extending previous works, we also study photo-realistic image enhancement using the MIT5K DSLR dataset [5]. We use 1000 images for training, and the standard split of 500 images for testing (as in [75]).
Finally, as previous works [43, 102, 62], we combine all the aforementioned training datasets, and we train our unified model for all-in-one restoration. Note that we do not include more real-world datasets because previous works do not provide results (or models) for those. Moreover, previous works were limited to synthetic data, in contrast, InstructIR also tackles real-world image enhancement.
Table 2: Quantitative results on five restoration tasks (5D) with state-of-the-art general image restoration and all-in-one methods. We highlight the reference model without text (image only), the best overall results, and the second best results. We also present the ablation study of our multi-task variants (from 5 to 7 tasks — 5D, 6D, 7D). This table is based on Zhang et al. IDR [102].
@@
Table 3: Comparisons of all-in-one restoration models for 3 restoration tasks (3D). We also show an ablation study for image denoising -the fundamental inverse problem- considering different noise levels. We report PSNR/SSIM metrics. Table based on [62].
@@
InstructIR w/o text 26.84/0.948 34.02/0.960 33.70/0.929 30.94/0.882 27.78/0.780 30.65/0.900
4.3 Multiple Degradation Results
We define two initial setups for multi-task restoration:
In Table 2, we show the performance of 5D models. Following Zhang et al. [102], we compare InstructIR with several state-of-the-art methods for general image restoration [95, 9, 10, 45, 96], and all-in-one image restoration methods [102, 43, 76, 21, 46]. We can observe that our simple image model (just 16M parameters) can tackle successfully at least five different tasks thanks to the instruction-based guidance and achieves the most competitive results. In Table 3 we can appreciate a similar behavior, when the number of tasks is just three (3D), our model improves further in terms of reconstruction performance.
Table 4: Ablation study on the sensitivity of instructions. We report PSNR/SSIM metrics for each task using our 5D base model. We repeat the evaluation on each test set 10 times, each time we sample different prompts for each image, and we report the average results. The “Real Users $\dagger$" in this study are amateur photographers, thus, the instructions were very precise.
Language Level Deraining Denoising Deblurring LOL
Basic & Precise 36.84/0.973 31.40/0.887 29.47/0.887 23.00/0.836
Basic & Ambiguous 36.24/0.970 31.35/0.887 29.21/0.885 21.85/0.827
Test Instructions.
InstructIR requires as input the degraded image and the human-written instruction. Therefore, we also prepare a test set of prompts i.e. instruction-image test pairs. The performance of InstructIR depends on the ambiguity and precision of the instruction. We provide the ablation study in Table 4. InstructIR is quite robust to more/less detailed instructions. However, it is still limited with ambiguous instructions such as “enhance this image". We show diverse instructions in Figures 5 and 6.
5 Multi-Task Discussion and Study
How does 6D work?
Besides the 5 basic tasks -as previous works-, we include single image super-resolution (SISR). For this, we include as training data the DIV2K [1]. Since our model does not perform upsampling, we use the Bicubic degradation model [1, 15] for generating the low-resolution images (LR), and the upsampled versions (HR) that are fed into our model to enhance them. Adding this extra task increases the performance on deblurring –a related degradation–, without harming notably the performance on the other tasks.
Table 5: Real Image Enhancement results on MIT5K [5].